A capacitor C1 of capacitance 1microF and a capacitor of capacitance 2microF are separately charged by a common battery for a long time. The two capacitors are then separately charged through equal resistors. Both the discharge circuits are connected at t=0. Then (multiple choice qstn)
(a) the current in each of the two discharging circuits is 0 at t=0
(b) the currents in the two discharging circuits at t=0 are equal but not 0
(c) the currents in the two discharging circuits at t=0 are unequal
(d) C1 loses 50% of its initial charge sooner than C2 loses 50% of its initial charge

Grade:12

2 Answers

Nirmal Singh.
10 years ago
As in case of discharging of a capacitor though a resistance
q = q0*e^-t/CR
i = -dq/dt = (q0/CR)*e^-t/CR
at t = 0, i = q0 / CR
we know q0 = CV
so i = CV / CR
i = V / R
i.e. the current at t = 0 is independent of capacitance in the circuit and as here V and R is same for both the circuit, the current at t = 0 will be same [t V/R] and maximum in both the circuits.
In case of discharging of a capacitor through a resistance as
q = q0*e^-t/CR
t = CR loge [q0 / q]
So time taken for 50% of the charge to leak through the resistance


i.e. the capacitor C1 will be lose its 50% charge sooner than CO2
Hence option (b) and (d) are correct
Thanks & Regards,
Nirmal Singh
Askiitians Faculty
vaanya
10 years ago
thanku :)